40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Entry  Wed Apr 20 00:12:12 2022, Yehonathan, Update, WOPO, Still figuring out the readout electronics and fixing of some stuff 20220419_153109.jpg20220419_164600.jpg20220419_184248.jpg20220419_184233.jpg
    Reply  Thu Apr 21 10:33:33 2022, shruti, Update, WOPO, Still figuring out the readout electronics and fixing of some stuff 
       Reply  Mon Apr 25 11:08:35 2022, Yehonathan, Update, WOPO, Still figuring out the readout electronics and fixing of some stuff 
          Reply  Thu Apr 28 14:12:20 2022, Yehonathan, Update, WOPO, Still figuring out the readout electronics and fixing of some stuff JustTIR.pdfTIR_WithCapacitors.pdfOnlyV_plus.pdf
Message ID: 2759     Entry time: Wed Apr 20 00:12:12 2022     Reply to this: 2760
Author: Yehonathan 
Type: Update 
Category: WOPO 
Subject: Still figuring out the readout electronics and fixing of some stuff 

{Yehonathan, Shruti}

1. Grabbed 30Hz-3GHz HP spectrum analyzer from the Cryolab. Installed it in the WOPO lab under the optical table. We figured out how to do a zero-span measurement around 10MHz. The SA has only one input so we try to combine the signals with an RF splitter. We test this capability by sourcing the RF splitter with 10MHz 4Vpp sine waves from a function generator and measuring the output with a scope. We measure with the scope 1.44Vpp for each channel. The combined channel was 2.73Vpp. We then realized that we still don't have a way to adjust the gains electronically, so we moved on to trying the RF amplifiers (ZFL500 LN).

We assemble two amps on the two sides of a metal heatsink. We solder their DC inputs such that they are powered with the same wire (Attachment 1). We attach the heatsink to the optical table with an L bracket (Attachment 2).

We powered the amps using a 15V DC power supply and tested them by feeding them with 10MHz 10mVpp sine waves from a function generator. We observe on a scope an amplification by a factor of ~ 22. Which makes a power amplification of ~ 26db consistent with the amplifiers' datasheet.

We couldn't find highpass filters with a cutoff around 1MHz, so we resumed using the DC blocks, we test them by feeding white noise into them with a function generator and observing the resulting spectrum. First, we try the DC blocks with a 50 Ohm resistor in parallel. That happened to just cut the power by half. We ditch the resistor and get almost unity transmission above 20kHz.

Moving on to observing LO shot noise, we open the laser shutter. We find there is only 0.7mW coming out of each port of the fiber BHD BS. We measure the power going into the BS to be 4mW. This means the coupling between the LO fiber and the BS fiber is bad. We inspect the fibers and find a big piece of junk on the BS fiber core. We also find a small particle on the LO fiber side. We cleaned both fibers and after butt coupling them we measure 1.6mW at each port. We raise this power to 2mW per port.

We connect the outputs of the PDs to the amps through the DC blocks. The outputs of the amps were connected to the Moku's inputs. The PDs were responding very badly and their noise was also bad. We bypass the amps to debug what is going on. We connect the PDs to a scope. We see they have 300mV (attachment 3) dark noise which is super bad and that they hardly respond to the light impinging on them (attachment 4). We shall investigate tomorrow.

Attachment 1: 20220419_153109.jpg  409 kB  Uploaded Wed Apr 20 16:29:02 2022  | Hide | Hide all
20220419_153109.jpg
Attachment 2: 20220419_164600.jpg  681 kB  Uploaded Wed Apr 20 16:29:26 2022  | Hide | Hide all
20220419_164600.jpg
Attachment 3: 20220419_184248.jpg  549 kB  Uploaded Wed Apr 20 16:29:37 2022  | Hide | Hide all
20220419_184248.jpg
Attachment 4: 20220419_184233.jpg  423 kB  Uploaded Wed Apr 20 16:29:49 2022  | Hide | Hide all
20220419_184233.jpg
ELOG V3.1.3-